OpenStudy (samsan9):

A local store owner pays her employees time-and-a-half for overtime. That means for every hour an employee works more than 40 hours per week, the store will pay 1.5 times the regular hourly wage of $8.50. Combine the two functions you just wrote to form a piecewise function, W(x), in which each function is defined for a specific domain.

OpenStudy (samsan9):

sorry there was another part to this problem that was solved earlier

OpenStudy (loser66):

Your attempt?

OpenStudy (samsan9):

i dont really know what a piecewise fnction is...

OpenStudy (samsan9):

260+(h-40)(9.75) do you think this could be the answer?

OpenStudy (loser66):

Now, just regular working ( I mean not over time) the salary is S1 = 8.50*h , h \(\leq \)40 where h is hour

OpenStudy (loser66):

Over time function S2= 340 + 12.75 *(h-40) when h >40 Let me explain why we have those numbers 340 =8.5*40 because for 40 hours, they are paid regular salary 12.75 = 1.5*8.5, this number is overtime salary. This salary applied for overtime only. So that it is timed to the hours they work more than usual. right? for example, you work 41 hours. the first 40 hours costs 40*8.5 = 340 the 1more hour cost 12.75 *1 = 12.75 total, you get 352.75 , right? Now, apply our formula S2= 340 +12.75 * (41-40) =352.75 THE SAME

OpenStudy (samsan9):

can you help me with the last part?

OpenStudy (samsan9):

Determine how much an employee will be paid if he or she were to work 50 hours in a week.

OpenStudy (loser66):

\[W(x)=\begin{cases}8.50*x ~~~~if~~~x\leq 40\\340+12.75(x-40)~~~if~~x>40\end{cases}\]

OpenStudy (samsan9):

is that the answer?